Determine the point of intersection of the following system of linear equations graphically.
y=-1/2x + 3
3y-2x=-12

Answer by mananth(12270)

0.5 x + 1 y = 3 .............1
-2 x + 3 y = 12 .............2

multiply (1)by -3
Multiply (2) by 1
-1.5 x -3 y = -9
-2 x 3 y = 12
Add the two equations
-3.5 x = 3
/ -3.5
x = -0.86 OR - 6/7
plug value of x in (1)
0.5 x + 1 y = 3
-0.43 + 1 y = 3
1 y = 3 + 0.43
1 y = 3.43
y = 3.43 OR 3 3/7